OpenStudy (lgbasallote):

In a poll of 500,000 people, if the occurrence of very wealthy people is 3 out of every 10,000, what is the probability that there exists between 150 to 152 (inclusive) very wealthy people among the 500,000?

OpenStudy (anonymous):

Can we use normal distribution ?

OpenStudy (lgbasallote):

i doubt....but let me see....

OpenStudy (lgbasallote):

^by that i meant show me your proposed solution...

OpenStudy (anonymous):

n=500,000 , p =3/10,000 , q= 1-p Now, Mean = np = 150 Standard deviation = (npq)^1/2 = 12.2456 Then, P(150<=X<=152) P((150-150)/12.2456 <=z<=(152-150)/12.2456) P(0<=z<=0.1633) P(z<=0.1633) - P(z<=0)

OpenStudy (lgbasallote):

hmm...i don't think that's applicable...

OpenStudy (lgbasallote):

this is actually a poisson distribution

OpenStudy (lgbasallote):

i suppose no one knows what it is.....